Operational Report 28-31OCT2021

MDRS Operations Report 28-31-OCT-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Bathroom leak is worse

Action taken for non-nominal systems: Will investigate as soon as possible

Generator: I attempted to put the generator back on autostart today and something went wrong and we lost power. I got power back online but will not attempt to put the generator online again until tomorrow. Power was out for about an hour Sunday from about 7-8 pm while I troubleshot the fault, which was VDC, meaning something went wrong with the programming in the generator.

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— SOC % Last 24 hours (Sunday):

Average 72.6

Minimum 54

Maximum 100

Notes on solar: November is the cut off for running on solar only.

Propane Reading, station tank – 64%

Propane Reading, director tank— 55%

Propane Reading, intern tank— 72+ %

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 1 gallons

Water (loft tank): 60 gallons

Water Meter: not checked units

Water (static tank) – 400 gallons

Static to Loft Pump used – yes

Water in GreenHab – 80 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 100 gallons

Hab toilet tank emptied: yes

Perseverance rover used: yes

Hours: 221.6

Beginning charge: 100

Ending charge: 22

Currently charging: yes

Sojourner rover used: yes

Hours: 155.4

Beginning charge: 100

Ending charge: 89

Currently charging: yes

Spirit rover used: yes

Hours: 142.0

Beginning charge: 100

Ending charge: 54

Currently charging: yes

Opportunity rover used: yes

Hours: 73.3

Beginning charge: 100

Ending charge: 88

Currently charging: yes

Curiosity rover used: no

Hours: 149.6

Beginning charge: 100

Ending charge: 70

Currently charging: yes

Notes on rovers: We took all five rovers on the same trip. Trip charges varied wildly:

Soji has new batteries and has been used consistently most days and has the best charge.

Opportunity has new batteries but has not been use much due to overheating, and it overheated at the halfway point of our trip.

Curi and Spirit have new batteries but they have not been used enough to fully charge them, but I’m still not such why they were so far apart from each other in remaining charge.

Percy is used a lot but does not have new batteries. This loss of charge is concerning as these batteries are only about at the halftime or less of their life.

NOTE FOR FUTURE OPERATIONS: If you try to charge all 5 rovers at once, they draw a lot of power. Don’t charge late in the afternoon if you are only using solar power.

ATV’s Used: (Honda, 350.1, 350.2, 350.3): One 350 used for short trip to Kissing Camel

Reason for use: Mission Planning

Oil Added? No

ATV Fuel Used: part of a gallons

# Hours the ATVs were Used today: 1

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town for water

CrewCar used and why, where? Yes, to town for food for crew

Luna used and why, where? Yes, To Grand Junction. Oil change was performed. Air filter needs replacement.

General notes and comments: Nothing to report

Summary of internet: All accounts are nominal.

Summary of suits and radios: Nothing to report. Headsets and cleaning clothes have been cleaned and put in individual packs to be given to each crewmember.

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: Nothing to report

Summary of Hab operations: Added alarm to the loft water tank, which goes off if the water level in the tank gets too high. You are welcome, 200 plus crews who complained about it, and especially to those crewmembers who had their staterooms flooded.

Summary of Outpost operations: Assistant Director is in residence and he has begun cleaning the intern trailer, which is a mess due to work that was done on it. Both residences are using wall heaters until we get the generator operational.

Summary of GreenHab operations: We harvested all the greens as they could not wait for the next crew. Replanted new greens. Currently heating at 70 degrees F.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Robotic Observatory is fully operational.

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Nothing at this time.

Copyright © The Mars Society. All rights reserved. | Main Site